A mathematics teacher in senior high school developed a problem-solving test to randomly selected 40 grade 11 

students. These students had an average score of 85 and a standard deviation of 5. If the population had a mean 

score of 90 and a standard deviation of 3, use 5% level of significance to test the hypothesis.


1
Expert's answer
2021-07-29T11:50:43-0400

The following null and alternative hypotheses need to be tested:

Null hypothesis: the population mean is equal to 90.

Alternative hypothesis: the population mean is not equal to 90.

"H_0:\mu=90"

"H_1:\mu\not=90"

The significance level is "\alpha=0.05."

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

The critical value for a two-tailed test is "z_c=z_{1-\alpha/2}=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{x-\mu}{\sigma/\sqrt{n}}=\dfrac{85-90}{3/\sqrt{40}}\approx-10.541"

Since it is observed that "|z|=10.541>1.96>z_c," is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=2P(Z<-10.541)=0," and since "p=0<0.05=\alpha," is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 90, at the "\alpha=0.05"  significance level.